# Complexity Bounds for Smooth Convex Multiobjective Optimization

### v1

@@ 5 Upper Bounds via Scalarization and Rate Comparison
    $x^{k+1}=y^{k}-\tfrac{1}{L}\,\nabla f_{\lambda}(y^{k}),\qquad t_{k+1}=\tfrac{1+\sqrt{1+4t_{k}^{\,2}}}{2},\qquad y^{k+1}=x^{k+1}+\tfrac{t_{k}-1}{t_{k+1}}\big{(}x^{k+1}-x^{k}\big{)}.$
@@
    $x^{k+1}=y^{k}-\tfrac{1}{L}\,\nabla f_{\lambda}(y^{k}),\qquad y^{k+1}=x^{k+1}+\beta\big{(}x^{k+1}-x^{k}\big{)}.$
@@
whence $\min_{0\leq t<T}\|\nabla f_{\lambda}(x^{(t)})\|\leq LR_{\lambda}/\sqrt{T}$. Since $\mathcal{G}(x)\leq\|\nabla f_{\lambda}(x)\|$, we obtain $\min_{0\leq t<T}\mathcal{G}(x^{(t)})\leq LR_{\lambda}/\sqrt{T}$ for the oblivious one-step class; this is tight in general for that class.
@@ 6 Conclusion
We developed an oracle complexity framework for smooth multiobjective optimization (MOO) focused on the task of finding $\varepsilon$-Pareto stationary points. Our lower bounds cover both convex and strongly convex regimes and distinguish between algorithm classes. In the $\mu$-strongly convex case, we proved a tight linear-rate lower bound that implies a worst-case iteration complexity of $\Theta\!\big{(}\sqrt{\kappa}\log(1/\varepsilon)\big{)}$ with $\kappa=L/\mu$, matching classical single-objective theory when specialized to scalarizations. In the merely convex case, we established an $\Omega(1/T)$ min-iterate lower bound for oblivious one-step methods and a universal $\Omega\!\big{(}1/T^{2}\big{)}$ last-iterate lower bound for fully adaptive span first-order methods, thereby clarifying the landscape across non-adaptive and adaptive algorithms.
On the positive side, we showed that applying an optimal scalar method (e.g., AGD) to a fixed scalarization attains an $O(1/T)$ upper bound for the Pareto stationarity gap; this matches our $\Omega(1/T)$ rate up to the difference in algorithm classes (oblivious span vs. oblivious one-step). We also calibrated upper and lower bounds via the initialization radii $R$ and $R_{\lambda}$ and proved that our constructions are robust: they apply to instances with genuinely distinct objectives and non-singleton Pareto fronts. Altogether, the results position MOO on a footing analogous to single-objective optimization while revealing new, vector-specific phenomena.
Two main directions remain open. First, in the convex setting for adaptive span methods, there is a gap between our universal $\Omega\!\big{(}1/T^{2}\big{)}$ last-iterate lower bound and known $O(1/T)$ upper bounds: closing this — either by stronger lower bounds (e.g., universal $\Omega(1/T)$) or by matching algorithms under the same oracle model — is a central challenge. Second, extending the framework to broader models (stochastic oracles, inexact gradients, constraints/composite objectives, and nonconvex MOO), as well as investigating higher-order information and dimension dependence, would further refine the fundamental limits that govern multiobjective first-order methods.
Appendix A Appendix: Derivations of Polynomial Extremal Results
@@ A.1 A tight $\Omega(1/T)$ lower bound for oblivious one-step methods (with $L$-capped steps)
Theorem A.1 (Tight $\Omega(1/T)$ product-form extremal under $L$-capped steps).
@@
Lemma A.3 (A product lower bound).
@@
$\prod_{i=1}^{r}(1-x_{i})\;=\;\Bigl{(}\prod_{i=1}^{r-1}(1-x_{i})\Bigr{)}(1-x_{r})\;\geq\;\Bigl{(}1-\sum_{i=1}^{r-1}x_{i}\Bigr{)}(1-x_{r})\;=\;1-\sum_{i=1}^{r}x_{i}\;+\;x_{r}\sum_{i=1}^{r-1}x_{i}\;\geq\;1-\sum_{i=1}^{r}x_{i},$
@@
Proof of Theorem A.1.
Fix $t\geq 1$ and a stepsize sequence $\alpha_{0},\ldots,\alpha_{t-1}$ with $0\leq\alpha_{k}\leq 1/L$ for all $k$. For any $\zeta\in[0,L]$, set $x_{k}:=\alpha_{k}\,\zeta\in[0,1]$. Applying Lemma A.3 gives
@@
$\max_{\zeta\in[0,L]}\Psi_{t}(\zeta)\;=\;\Psi_{t}(\zeta^{\diamond})\;=\;\frac{L}{t+1}\left(1-\frac{1}{t+1}\right)^{t}\;\leq\;\frac{L}{t+1}\,\mathrm{e}^{-\,t/(t+1)}\;\leq\;\frac{L}{\mathrm{e}\,(t+1)}.$
@@
We now translate Theorem A.1 into an oracle lower bound for the gradient norm.
Corollary A.4 (Oblivious one-step lower bound for the gradient norm).
@@
by Theorem A.1. ∎
This result is lifted to the multiobjective setting using the construction from Theorem 4.5. For the instance $f_{i}(x)=g(x_{V})+\frac{L}{2}\|x_{W}-a_{i}\|^{2}$, we have $\mathcal{G}(x)\geq\|\nabla g(x_{V})\|$ and $\mathrm{dist}(x^{(0)},\mathcal{P})=R$. Corollary A.4 thus directly implies a lower bound on $\min_{s\leq t}\mathcal{G}(x^{(s)})$. This establishes the bound for a non-degenerate MOO instance with distinct objectives and a non-singleton Pareto set.
A.2 Universal Markov-Based Bound for Span Methods
The lower bound for adaptive methods relies on $\inf_{p_{t}}\max_{\zeta\in[0,L]}|\zeta p_{t}(\zeta)|$ where $p_{t}$ is any polynomial with $\deg p_{t}\leq t$ and $p_{t}(0)=1$.
Lemma A.5 (Markov Brothers’ Inequality (see [28], Chapter 2, Section 2.7, Page 123)).
For any polynomial $q(x)$ of degree $n$, $\max_{x\in[-1,1]}|q^{\prime}(x)|\leq n^{2}\max_{x\in[-1,1]}|q(x)|$.
@@ A.3 Chebyshev Extremal Problem on $\mu,L]$
The lower bound for strongly convex methods requires solving $\min_{p_{t}}\max_{\zeta\in[\mu,L]}|p_{t}(\zeta)|$ subject to $\deg p_{t}\leq t$ and $p_{t}(0)=1$. The affine transformation $\xi=\frac{2\zeta-(L+\mu)}{L-\mu}$ maps the interval $\zeta\in[\mu,L]$ to $\xi\in[-1,1]$. The constraint $p_{t}(0)=1$ is evaluated at $\zeta=0$, which corresponds to $\xi_{0}=-\frac{L+\mu}{L-\mu}$. Let $\tilde{p}_{t}(\xi)=p_{t}(\zeta(\xi))$. The problem becomes:
@@
$\rho=\frac{\kappa+1}{\kappa-1}+\sqrt{\left(\frac{\kappa+1}{\kappa-1}\right)^{2}-1}=\frac{\kappa+1+2\sqrt{\kappa}}{\kappa-1}=\frac{(\sqrt{\kappa}+1)^{2}}{(\sqrt{\kappa}-1)(\sqrt{\kappa}+1)}=\frac{\sqrt{\kappa}+1}{\sqrt{\kappa}-1}.$

### v2

@@ 5 Upper Bounds via Scalarization and Rate Comparison
    $x^{k+1}$ $=y^{k}-\tfrac{1}{L}\,\nabla f_{\lambda}(y^{k}),$
    $t_{k+1}$ $=\tfrac{1+\sqrt{1+4t_{k}^{\,2}}}{2},$
    $y^{k+1}$ $=x^{k+1}+\tfrac{t_{k}-1}{t_{k+1}}\bigl(x^{k+1}-x^{k}\bigr).$
@@
    $x^{k+1}$ $=y^{k}-\tfrac{1}{L}\,\nabla f_{\lambda}(y^{k}),$
    $y^{k+1}$ $=x^{k+1}+\beta\bigl(x^{k+1}-x^{k}\bigr).$
@@
Remark 5.6 (Tighter scalar upper bounds via gradient-minimization methods).
The bounds in Theorem 5.4–Corollary 5.5 are obtained by applying AGD to a fixed scalarization $f_{\lambda}$ and then translating function-value convergence into a gradient-norm bound via Lemma 5.1 (using that $G(x)\leq\|\nabla f_{\lambda}(x)\|$). If one is specifically interested in gradient-norm complexity for the scalar problem $\min_{x}f_{\lambda}(x)$, then tighter results are available in the single-objective literature.
In particular, Lan, Ouyang and Zhang [22] develop optimal (and parameter-free) first-order methods for computing $\hat{x}$ with small (projected) gradient norm. Specializing to our unconstrained setting (where projected gradient reduces to the gradient), their results imply that for convex $L$-smooth $f_{\lambda}$ one can obtain $\|\nabla f_{\lambda}(\hat{x})\|\leq\varepsilon$ within $\mathcal{O}\!\left(\sqrt{L}\,R_{\lambda}/\varepsilon\right)$ gradient evaluations, where $R_{\lambda}=\|x^{(0)}-x^{\star}_{\lambda}\|$. In the $\mu$-strongly convex case, they obtain $\mathcal{O}\!\left(\sqrt{L/\mu}\,\log(\|\nabla f_{\lambda}(x^{(0)})\|/\varepsilon)\right)$ gradient evaluations. Combining these guarantees with $G(x)\leq\|\nabla f_{\lambda}(x)\|$ yields the corresponding Pareto-gap upper bounds for the multiobjective problem when optimizing a fixed scalarization.
@@
whence $\min_{0\leq t<T}\|\nabla f_{\lambda}(x^{(t)})\|\leq LR_{\lambda}/\sqrt{T}$. Since $\mathcal{G}(x)\leq\|\nabla f_{\lambda}(x)\|$, we obtain $\min_{0\leq t<T}\mathcal{G}(x^{(t)})\leq LR_{\lambda}/\sqrt{T}$ for the oblivious one-step class.
@@ 6 Conclusion
We developed an oracle-complexity theory for finding $\varepsilon$-Pareto stationary points in smooth multiobjective optimization, measured by the Pareto stationarity gap $\mathcal{G}(x)$. A central tool is a robust non-degenerate lifting construction that embeds scalar hard instances into MOO instances with distinct objectives and non-singleton Pareto fronts while preserving lower bounds on $\mathcal{G}$.
In the $\mu$-strongly convex regime we obtained a tight linear-rate lower bound for span methods matching accelerated upper bounds. In the convex regime we separated algorithmic power: oblivious one-step methods admit an $\Omega(1/T)$ min-iterate lower bound, while oblivious span methods satisfy a universal polynomial-degree $\Omega(1/T^{2})$ last-iterate bound. We then showed that the polynomial-degree limitation is loose: by lifting the geometric stationary-point lower bounds of Carmon–Duchi–Hinder–Sidford, we proved an $\Omega(\sqrt{L\Delta}/T)$ min-iterate lower bound for adaptive first-order methods under bounded scalarization gap, recovering $\Omega(LR/T)$ on our lifted family. Finally, leveraging that the lifting does not require convexity of the embedded function, we extended the framework to nonconvex MOO under Lipschitz gradients, obtaining the tight-in-order $\Omega(\sqrt{L\Delta/T})$ complexity.
Several directions remain open. Most prominently, the last-iterate complexity of adaptive first-order methods in convex MOO under bounded scalarization gap (or under geometric distance models) is not yet fully characterized. Extending the framework to randomized methods, stochastic oracles, constraints/composite objectives, and refined dimension-dependent regimes are further natural directions for a comprehensive complexity theory of Pareto stationarity.
Appendix A Derivations of Polynomial Extremal Results
@@ A.1 A tight $\Omega(1/T)$ lower bound for oblivious one-step methods (with $L$-capped steps)
Theorem A.1.1 (Tight $\Omega(1/T)$ product-form extremal under $L$-capped steps).
@@
Lemma A.1.3 (A product lower bound).
@@
$\prod_{i=1}^{r}(1-x_{i})$ $=\Bigl(\prod_{i=1}^{r-1}(1-x_{i})\Bigr)(1-x_{r})$
$\geq\Bigl(1-\sum_{i=1}^{r-1}x_{i}\Bigr)(1-x_{r})$
$=1-\sum_{i=1}^{r}x_{i}\;+\;x_{r}\sum_{i=1}^{r-1}x_{i}$
$\geq 1-\sum_{i=1}^{r}x_{i}.$
@@
Proof of Theorem A.1.1.
Fix $t\geq 1$ and a stepsize sequence $(\alpha_{k})_{k=0}^{t-1}$ with $0\leq\alpha_{k}\leq\tfrac{1}{L}$ for all $k$. For any $\zeta\in[0,L]$, set $x_{k}:=\alpha_{k}\,\zeta\in[0,1]$. Applying Lemma A.1.3 gives
@@
$\max_{\zeta\in[0,L]}\Psi_{t}(\zeta)$ $=\Psi_{t}(\zeta^{\diamond})$
$=\frac{L}{t+1}\left(1-\frac{1}{t+1}\right)^{t}$
$\leq\frac{L}{t+1}\,\mathrm{e}^{-\,t/(t+1)}$
$\leq\frac{L}{\mathrm{e}\,(t+1)}.$
@@
We now translate Theorem A.1.1 into an oracle lower bound for the gradient norm.
Corollary A.1.4 (Oblivious one-step lower bound for the gradient norm).
@@
by Theorem A.1.1. ∎
This result is lifted to the multiobjective setting using Theorem 4.1: for the instance $f_{i}(x)=g(x_{V})+\frac{L}{2}\|x_{W}-a_{i}\|^{2}$, we have $\mathcal{G}(x)\geq\|\nabla g(x_{V})\|$ and $\mathrm{dist}(x^{(0)},\mathcal{P})=R$. Corollary A.1.4 thus directly implies a lower bound on $\min_{s\leq t}\mathcal{G}(x^{(s)})$. This establishes the bound for a non-degenerate MOO instance with distinct objectives and a non-singleton Pareto set.
A.2 Universal Markov-Based Bound for Oblivious Span Methods
The lower bound for oblivious methods relies on $\inf_{p_{t}}\max_{\zeta\in[0,L]}|\zeta p_{t}(\zeta)|$ where $p_{t}$ is any polynomial with $\deg p_{t}\leq t$ and $p_{t}(0)=1$.
Lemma A.2.1 (Markov Brothers’ Inequality (see [32], Chapter 2, Section 2.7, Page 123)).
For any polynomial $q(x)$ of degree $n$,
$\max_{x\in[-1,1]}|q^{\prime}(x)|\leq n^{2}\max_{x\in[-1,1]}|q(x)|.$
@@ A.3 Chebyshev Extremal Problem on $\mu,L]$
The lower bound for strongly convex methods requires solving
$\min_{p_{t}}\;\max_{\zeta\in[\mu,L]}\lvert p_{t}(\zeta)\rvert\quad\text{s.t.}\quad\deg p_{t}\leq t,\;p_{t}(0)=1.$
The affine transformation $\xi=\frac{2\zeta-(L+\mu)}{L-\mu}$ maps the interval $\zeta\in[\mu,L]$ to $\xi\in[-1,1]$. The constraint $p_{t}(0)=1$ is evaluated at $\zeta=0$, which corresponds to $\xi_{0}=-\frac{L+\mu}{L-\mu}$. Let $\tilde{p}_{t}(\xi)=p_{t}(\zeta(\xi))$. The problem becomes:
@@
$\rho$ $=\frac{\kappa+1}{\kappa-1}+\sqrt{\left(\frac{\kappa+1}{\kappa-1}\right)^{2}-1}$
$=\frac{\kappa+1+2\sqrt{\kappa}}{\kappa-1}$
$=\frac{(\sqrt{\kappa}+1)^{2}}{(\sqrt{\kappa}-1)(\sqrt{\kappa}+1)}$
$=\frac{\sqrt{\kappa}+1}{\sqrt{\kappa}-1}.$